--++ vidyo

vidyo link

Action Items:

News:

  • welcome to Andres Tico, new post-doc since June 21st. Presentation from him next Thursday.
  • Cristina will finish her DT ... suggestions about new candidates are welcome.
  • Andrea Triossi is now in ciemat. There was a strong interest to have him in Padova, the director did not agreed about conditions for him to stay.
  • RUN STATUS:
    • Marina - re-start again - 3 runs, one with 246 bunches and PU40. Now pixel threshold scan. Next fill would be with full machine. after this we should start with full bunch data-taking and full lumi.
    • first layer of pixel. ROC seems not to work as expected and solution is not clear up to now.

Framework e produzione

Attivita' di analisi

B physics

  • good news - we managed to include two paths for JPsi phi
  • bad news - Alessio computed the rates of old paths. Overall is fine but we have an inefficiency in the tracking. Once it is fixed, the increment of factor 3-4 for the rate is expected. need to find some people to re-study bsmumu path (does Nicola P. have time?).
  • bad news - P5' presentation at conference --> questions from LHCb people. Possible issue related to bkg prediction which would lead to a bias. It should be checked/fixed soon, before CWR. Best before Stefano's talk at EPS.

Dark Matter

  • Siew Yan has won the phD position here in Padova. Cariparo fellowship. He will come Oct 1st. From Fermilab they asked him to stay a bit longer. He is getting experience on big data tools.

Dibosoni, VZ risonante (V=W,Z)

  • good news -> ARCs gave green-light. In track for EPS -- * Approval tomorrow at 2pm *
  • Poster from Lisa at EPS.
  • Paul is certified language editor of it.

hh4b non risonante

  • pre-approval got two weeks ago.
  • further checks made and two steps of ARCs questions replied.
  • unblinding performed in agreement with ARCs --> fluctuation found in data. Surely due to background prediction. Bias needed to be reviewed/improved.
  • Andre David said we are not pre-approved... back to step zero?
  • out from EPS timeline.

  • tracker TDR finished CWR. Section on hh4b is included (study on signal).

Same sign leptons

Seesaw a 13 TeV

EPR

  • Pablo is involved in the development in the new tagger. Work needed within August.

Trigger

Detector

  • Lisa is doing all the 'burocracy' needed to start working at GIF for DT.

  • from Legnaro:
    • bicells loose 20% gain after a radiation larger than expected from HL-LHC
    • MB1 loose 70% with just 1/10 of radiation expected from HL-LHC.
    • Unique difference is the ...

A.O.B.:

TESI: Lisa started writing, Alessio not yet. It will be their priority starting from next week.

-- FrancoSimonetto - 2017-06-27

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2017-06-27 - MartinoDallOsso
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback